OpenStudy (anonymous):

Write a polynomial function f of least degree that has rational coefficients, a leading coefficient of 1, and the zeros 2, -3, -3i

OpenStudy (anonymous):

start with \[(x-2)(x+3)(x+3i)(x-3i)\] and multiply out i would do \[(x+3i)(x-3i)=x^2+9\] first then multiply by the rest

OpenStudy (anonymous):

Wait why is there an equal sign right there???

OpenStudy (anonymous):

are you asking why \((x+3i)(x-3i)=x^2+9\)

OpenStudy (anonymous):

Yes, wouldn't you just multiply them all?

OpenStudy (anonymous):

you have to multiply them all, yes but what i meant was i would do that product first because it is the easiest

OpenStudy (anonymous):

\((s+3i)(x-3i)=x^2+3ix-3ix-9i^2=x^2+9\)

OpenStudy (anonymous):

Oooohh, I get it! Thanks so much!!